I can not figure out the answer to this word problem, one number is 3 less than 6 times the second number their sum is 46. find the numbers, could you please help me.

Answer by macston(5194)

x and y are the numbers
x=6y-3
x+y=46 Substitute for x.
6y-3+y=46
7y=49
y=7 ANSWER 1: One of the numbers is 7.
x+y=46
x=46-y
x=46-7=39 ANSWER 2: The other number is 39.
CHECK:
x=6y-3
39=6(7)-3
39=42-3
39=39